I need help with this question
I am a rectangle my perimeter is 60ft my lenght is twice as long as my width how much area do I cover?

Answer by ramkikk66(644)

Question:
I am a rectangle my perimeter is 60ft my lenght is twice as long as my width how much area do I cover?
Answer:

For a rectangle with width W and length L,
Perimeter = 2*(L+W)
Area = L*W

Let the width be W ft
Then length = 2*W ft
Perimeter = 2*(length + width) = 2*(W + 2*W) = 2*3*W = 6*W = 60 ft
Solving for W, W = 60/6 = 10 Ft and Length = 2*10 = 20 Ft.

Hence, area = length * width = 20*10 = 200 Sq feet.

Hope this helps!